Operation Summary for September 14, 2009 - September 16, 2009 |
Monday September 14 The midnight shift began with Operations monitoring the stash (127.2E10), with the Antiproton Source (Pbar) stacking (136.0E10), and with startup operations and studies in progress for the Tevatron (TeV), the Main Injector (MI), Recycler, NuMI, and MiniBooNE. The day shift started with MI-65 in access and with TeV ground fault detector repairs. TeV studies followed. HEP shot setup began and was promptly delayed by Lambertson (LAM62) power supply regulation problems. Antiprotons were transferred to the Recycler when appropriate. Pbar, NuMI, and MiniBooNE took beam when it was available. The evening shift began with a Klystron RF Debuncher water skid failure. Linac personnel investigated. The LAM62 power supply regulation problems were resolved by 7:30 PM. HEP shot setup had resumed by 9:30 PM and continued through to the end of shift. Pbar and MiniBooNE took beam when it was available. Tuesday September 15 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash and stack, and with NuMI and MiniBooNE taking beam. During the midnight shift, efforts to tune up reverse proton injection from the TeV to MI and ultimately extract antiprotons out of MI were largely unsuccessful. As a result, the TeV continued with proton only studies after ~90% of the available antiprotons were extracted in attempts to tune up the A150 beamline. Booster beam was turned off at 6 AM in preparation for an access. Operations also prepared MI and Recycler for accesses. Pbar and MiniBooNE took beam when available. The day shift started with Pbar cryo system repairs, which inhibited stacking and continued all shift. Booster, NuMI, and Main Injector personnel accessed during this period. TeV tuned with stored protons. The Booster access was completed by 10:30 AM and Booster was turned on. NuMI and MI were taking beam by 1 PM. TeV personnel looked at reverse injection beam and found a power supply (I:V901) that wasn’t regulating properly. EE Support personnel worked on V901 while NuMI personnel conducted target scans. The stack was transferred to the Recycler. Operators searched and secured MI-31. At 3:45 PM, the TeV ramp dumped while experts were verifying V901. TeV personnel and operators investigated the TeV problems and found a bad card in the Safety Coordinator. During the evening shift, the TeV A3 Safety Coordinator was repaired by 4:48 PM. NuMI studies ended and the NuMI Absorber Hall was in access by 5:32 PM; NuMI resumed taking beam at about an hour later. Operations put the TeV into shot setup at 6:02 PM, but the first pulse of beam quenched the TeV at sector F1 due to a kicker (F17K3) being off. The TeV had recovered by 7:53 PM. Shot setup resumed and store 7183 was colliding by 10:05 PM with an average initial luminosity (AIL) of 63.61e30. Operators prepared MI for access at 10:23 PM to repair an Electron Cooling vacuum valve. Work on the NuMI kicker fluorinert system was complete by 11:08 PM. Operators were turning on MI by shift’s end. MiniBooNE took beam when it was available. Short downtimes during the shift were due to problems with LRF2, KRF1, L:QPS302, and a MI-40 vacuum burst. Wednesday September 16 The midnight shift began with Operations monitoring store 7183 (44.41E30) and the stack (6.8E10), with the Recycler off due to the MI access, and with MI being brought online after an access. During the midnight shift, the MI turn on was completed and Operations established beam to NuMI and MiniBooNE by 12:30 AM. Store 7183 collided all shift and the Pbar cryo system repairs and cooldown continued. A TeV RF station (TRF2) tripped off with only a small impact on colliding beams. The NuMI horn tripped off towards the end of the shift. Operations contacted EE Support personnel to investigate. NuMI and MiniBooNE took beam when it was available. The Plans for Wednesday Due to the Pbar cryo system work, Pbar won’t be able to stack until around midnight. Recycler will conduct proton studies and then accept a small antiproton stash. Pbar will conduct a destructive study with what antiprotons are left after the stash. MI will commission their collimators. The TeV will keep the store in until noon or so, then conduct an end-of-store study, and then follow with a proton only study. |
|
More Information Comments and Suggestions |
| last modified 09/16/2009 email Fermilab |
| Security, Privacy, Legal | |